Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 10:33 Wednesday 21 November 2018 (7066) Print this report
About the BS condition

With Mark, Kozu-kun, Enzo and Hirata-san

Kozu-kun pointed out that the oplev was out of range. We checked the positions of the different fiters and they are as follows

  • F0: -6863 µm
  • F1: 243 µm
  • BF: -1325 µm (likely rendering the IM-V OSEMs out of range)

The displacement of F0 is particularly large. We took a look to the slow channel measurements. Several of the turning points still have to associated with a cause but now we have the following information:

  1. The plots show the minute average 100 days back from the 2nd of November.
  2. Yokozawa-san finished cabling o the 3rd f August (5709).
  3. We shut down the data acquisition system on Friday 7th of September and turn them back on on Tuesday 11th. There was a blackout on Monday 10th.
  4. F0 settles down aproximately to its current position on the 19th of October.
  5. From the 11th of September the F0 and F1 follow a similar trend.
  6. Temperature is shown but dates are yet to be written in the plot.
Images attached to this report
Comments to this report:
fabian.arellano - 15:12 Wednesday 21 November 2018 (7073) Print this report

After a correction in the calibration factor of the F0 LVDT (see entry 7066), the F0 LVDT readout is -2796 µm.

Search Help